### v3

# Engineering subgap states in superconductors by altermagnetism

Abstract
We investigate the realization and control of subgap states by tailored altermagnetic fields on unconventional superconductors. When the symmetries of altermagnetism and unconventional superconductivity align, we demonstrate the emergence of bulk zero-energy flat bands, giving rise to a zero-bias conductance peak. The symmetry and strength of $d$- and $g$-wave altermagnets strongly affect the surface Andreev states from $d$-wave and chiral $d$- and $p$-wave superconductors. As a result, distinct types of subgap states are realized, including curved and flat bands, that can be detected by tunneling spectroscopy. Furthermore, we find that the altermagnetism-induced subgap states give rise to a large spin conductance at zero net magnetization which helps identify the strength of the underlying altermagnetism and superconductivity. Our results offer a solid route for designing and manipulating subgap states in superconducting systems, which can be useful for functionalizing superconducting spintronic devices.
Introduction.— The ability to induce and control subgap states in superconducting systems is a central topic in condensed matter physics, with a profound fundamental relevance and high potential for quantum applications. One example of such subgap states is the surface Andreev bound states (SABSs) [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13], which emerge due to the anisotropic nature of the superconducting pair potentials and can be detected by tunneling spectroscopy [6]. Interestingly, SABSs in unconventional superconductors (SCs) can lead to Majorana states [14, 15, 16, 12], which hold promise for designing topological quantum bits [17, 18, 19, 20, 21, 22]. Moreover, certain superconducting setups can also host bulk subgap states, such as Bogoliubov-Fermi surfaces (BFSs) [23, 24, 25, 26, 27, 28, 29, 30, 31, 32], which signal the nontrivial topological nature of superconductivity [33, 27]. For practical purposes, subgap states need to be manipulated, often carried out by coupling external fields to their intrinsic quantum numbers [34, 35, 36], such as orbital or spin. Although this can indeed be achieved by ferromagnets [37, 38, 39], their unavoidable finite magnetization and stray fields make them detrimental for superconductivity [40], and therefore, other materials are required.
Figure 1: An altermagnet (AM, orange) coupled to a superconductor (SC, green) is used to engineer subgap states. A normal metal probe (N, gray) coupled to the AM-SC system at $x=0$ is then used to probe the induced subgap states via differential conductance $dI/dV$. (b) Table showing the types of emergent subgap states with respect to $k_{y}$ induced by altermagnetism in distinct unconventional SCs.
With the recent advent of altermagnetism [41, 42], a powerful avenue has emerged for coupling spins in superconductors thanks to the altermagnetic spin-orbit properties [43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53]. Altermagnets (AMs) are spin-compensated magnetic materials with a nonrelativistic momentum-dependent spin splitting and anisotropic spin-polarized Fermi surfaces [54], which result in zero net magnetization and hence vanishing stray fields unlike ferromagnets [55, 56]. In combination with SCs, AMs have already been shown to host intriguing properties [53], such as orientation-dependent Andreev reflections [57, 58, 59, 60, 61, 62], anomalous Josephson currents [63, 64, 65, 66, 67, 68, 69], nontrivial light-matter coupling [70, 71], superconducting diodes [72, 73, 74], spin-triplet Cooper pairs [75, 76, 77, 78, 79], Majorana states [80, 81, 82, 83, 84, 85], among other phenomena [86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96]. Despite the impressive advances, the use of altermagnetism for realizing subgap states is still an open problem.
In this work, we demonstrate the emergence and control of highly tunable subgap states by using altermagnetism in unconventional SCs. To show this, we focus on a hybrid system composed of a two-dimensional SC with altermagnetism due to its proximity to an AM [Fig. 1(a)], where a normal probe enables the detection of bulk and surface subgap states. We consider various types of unconventional SCs combined with $d$- and $g$-wave AMs, which allows us to reveal various types of subgap states, including bulk zero energy flat bands and SABSs and listed in Fig. 1(b). We further demonstrate that these subgap states give rise to strong charge and spin conductance signatures, enabling their detection and also the identification of the type of altermagnetism supporting them. Our findings demonstrate that altermagnetism can be effectively harnessed to generate tunable subgap states and spin currents in superconducting systems.
Model.— We aim to harness altermagnetism to realize and detect subgap states in unconventional SCs. We thus consider a junction along $x$, where an unconventional SC coupled to a two-dimensional AM is connected to a normal metal lead N at $x=0$, see Fig. 1(a). The AM induces an altermagnetic order parameter into the SC via proximity effect and the lead can probe surface and bulk properties. The altermagnetic superconducting junction is modeled by $\mathcal{H}=(1/2)\sum\nolimits_{\bm{k}}\hat{c}_{\bm{k}}^{{\dagger}}\mathcal{H}_{\bm{k}}\hat{c}_{\bm{k}}$, where the Nambu spinor is given by $\hat{c}_{\bm{k}}=(c_{\bm{k}\uparrow},c_{\bm{k}\downarrow},c_{-\bm{k}\uparrow}^{{\dagger}},c_{-\bm{k}\downarrow}^{{\dagger}})^{T}$, $\bm{k}=(k_{x},k_{y})$, and
$\mathcal{H}_{\bm{k}}=\left[\xi_{\bm{k}}+U(x)\right]\hat{\tau}_{z}+\mathcal{M}_{\bm{k}}(x)\hat{s}_{z}\hat{\tau}_{z}+\Delta_{\bm{k}}(x).$ (1)
Here, $\xi_{\bm{k}}=\hbar^{2}|\bm{k}|^{2}/(2m)-\mu$ is the kinetic energy, with $\mu$ being the chemical potential, $U(x)=U_{0}\delta_{x}$ is the interface barrier potential at $x=0$, with $U_{0}$ the barrier strength defining the parameter $Z=mU_{0}/(\hbar^{2}k_{F})$ that characterizes the tunneling ($Z\gg 1$) and full transparent ($Z=0$) regimes, $k_{F}=\sqrt{2m\mu/\hbar^{2}}$, and $\hat{s}_{i}(\hat{\tau}_{i})$ is the $i$-th Pauli matrix in spin (Nambu) space. Moreover, $\mathcal{M}_{\bm{k}}(x)=\Theta(x)h_{0}m_{\bm{k}}$ is the altermagnetic field, with $h_{0}$ denoting the altermagnetic strength and $m_{\bm{k}}$ the anisotropic factor, and, without loss of generality, the Neel vector of the AM is chosen along the $z$-axis. The unconventional superconducting pair potential $\Delta_{\bm{k}}(x)=\Theta(x)\hat{\Delta}_{\bm{k}}$ is only finite in the SC. We assumed $h_{0}\ll\mu$ so that the wavevectors at low energies can be approximated by $\bm{k}=k_{F}\bm{\hat{k}}$, with the unit vector $\bm{\hat{k}}=(\hat{k}_{x},\hat{k}_{y})$.
[…]
The quasiparticle energy dispersion in the SC with altermagnetism is given by
$E_{\bm{k}\nu}^{\pm}=\nu h_{0}m_{\bm{k}}\pm\sqrt{\xi_{\bm{k}}^{2}+|\Delta_{0}s_{\bm{k}}|^{2}}\,$, (2)
where $\nu=\pm$ indicating two possible anisotropic energy shifts ($\pm h_{0}m_{\bm{k}}$) of subbands. From Eq. (2), it is easy to see that, for $\nu h_{0}m_{\bm{k}}=|\Delta_{0}s_{\bm{k}}|$, we get $E^{+}_{\bm{k}\nu}\neq 0$ and $E^{-}_{\bm{k}\nu}=0$, the latter signaling a zero energy bulk flat band akin to BFSs [23]. This is an example of a bulk subgap state induced by altermagnetism, and its detection can be carried out via conductance in the setup of Fig. 1(a). Notably, this setup also allows us to search for altermagnetism-induced subgap states at the interface surface at $x=0$, which can be directly accessed via conductance as we explain below.
To obtain the conductance, we employ the scattering approach, which is based on the wavefunctions of injected electrons with spin $\sigma=\uparrow,\downarrow$. Due to translational invariance along $y$, the wavefunctions are $\Psi_{\sigma}(x,k_{y})=\psi_{\sigma}(x)e^{iyk_{y}}$. In N, $\psi_{\sigma}(x)$ is given by
$\psi_{\sigma}(x<0)=[\hat{A}_{\sigma}+a_{\sigma}\hat{B}_{\sigma}]e^{ixk_{x}}+b_{\sigma}\hat{A}_{\sigma}e^{-ixk_{x}}\,$, (3)
where $\hat{A}_{\uparrow}=\left(1,0,0,0\right)^{T}$, $\hat{A}_{\downarrow}=\left(0,1,0,0\right)^{T}$, $\hat{B}_{\downarrow}=\left(0,0,1,0\right)^{T}$ and $\hat{B}_{\uparrow}=\left(0,0,0,1\right)^{T}$, $k_{x}=k_{F}\cos\theta$, and $k_{y}=k_{F}\sin\theta$. The wave functions in the SC can be written in a similar way. Then, the coefficients $a_{\sigma}$ and $b_{\sigma}$ are obtained from the boundary conditions $\psi(0^{+})=\psi(0^{-})$, $\partial_{x}\psi(0^{+})-\partial_{x}\psi(0^{-})=(2mU_{0}/\hbar^{2})\psi(0)$. The angle-resolved charge (spin) conductance is found as [98, 99, 100] $\bar{\sigma}_{\sigma}^{e(s)}=1\pm|a_{\sigma}|^{2}-|b_{\sigma}|^{2}$, which gives
$\bar{\sigma}_{\sigma}^{e(s)}=\bar{\sigma}_{n}\frac{1\pm\bar{\sigma}_{n}|\Gamma_{\sigma}^{+}|^{2}+(\bar{\sigma}_{n}-1)|\Gamma_{\sigma}^{+}\Gamma_{\sigma}^{-}|^{2}}{|\Lambda-\eta(1-\bar{\sigma}_{n})\Gamma_{\sigma}^{+}\Gamma_{\sigma}^{-}|^{2}}\,$, (4)
with $\bar{\sigma}_{n}=1/[1+Z^{2}/(\cos^{2}\theta)]$. Here, $\Gamma_{\sigma}^{\pm}$ are given by
$\Gamma_{\sigma}^{\pm}=\frac{\Delta_{0}s_{\bm{k}_{\pm}}}{E_{\sigma,\bm{k}_{\pm}}+\sqrt{E_{\sigma,\bm{k}_{\pm}}^{2}-\Delta_{0}^{2}s_{\bm{k}_{\pm}}^{2}}}.$ (5)
where $E_{\uparrow,\bm{k}_{\pm}}=E-\mathcal{M}_{\bm{k}_{\pm}}$ and $E_{\downarrow,\bm{k}_{\pm}}=E+\mathcal{M}_{\bm{k}_{\pm}}$, with $\bm{k}_{+}$, $\bm{k}_{-}$ are given by $\bm{k}_{\pm}=k_{F}\left(\pm\cos\theta,\sin\theta\right)$. Moreover, $\Lambda=1$ for $d_{x^{2}-y^{2}}$- and $d_{xy}$-wave SCs; $\Lambda=e^{2i\theta}$ for chiral $p$-wave SCs; $\Lambda=e^{4i\theta}$ for chiral $d$-wave SCs. Also, $\eta=1$ for spin-singlet SCs (cases i,ii,iv), while $\eta=-1$ for spin-triplet SCs (case iii). Hence, the charge (spin) conductance in the superconducting state is calculated as
$\sigma_{\rm SC}^{e(s)}=\frac{e^{2}}{h}\int_{-\frac{\pi}{2}}^{\frac{\pi}{2}}\bar{\sigma}^{e(s)}\cos\theta d\theta.$ (6)
where $\bar{\sigma}^{e}=\bar{\sigma}_{\uparrow}^{e}+\bar{\sigma}_{\downarrow}^{e}$ and $\bar{\sigma}^{s}=\bar{\sigma}_{\uparrow}^{s}-\bar{\sigma}_{\downarrow}^{s}$ are the angle-resolved charge and spin conductances, respectively. $\sigma_{\rm SC}^{s}$ measures the ability to produce a spin current and plays a key role in superconducting spintronics. Later we normalize the conductances by $\sigma_{N}=(2e^{2}/h)\int\bar{\sigma}_{n}\cos\theta d\theta$.
At this point, it is important to highlight that, to access the surface ABSs states via conductance, the necessary condition is to set $\bar{\sigma}_{n}\rightarrow 0$ and find the zeroes of the denominator in Eq. (4), which leads to
$\Gamma_{\sigma}^{+}\Gamma_{\sigma}^{-}=\Lambda/\eta\,.$ (7)
This condition corresponds to the poles of the associated scattering matrix. Thus, the surface ABSs are obtained from Eq. (7). In what follows, we explore how the bulk subgap state discussed below Eq. (2) as well as surface ABSs can be induced by altermagnetism and detected by conductance in unconventional SCs. Our results in the continuum model are the same with the lattice model shown in the Supplementary Material [101].
Bulk flat bands by $d_{x^{2}-y^{2}}$-wave altermagnetism in $d_{x^{2}-y^{2}}$-wave SCs.— We begin by analyzing the junction formed by $d_{x^{2}-y^{2}}$-wave SC with $d$-wave altermagnetism. Before delving into the conductance, we remind our discussion below Eq. (2), where we show that at $\nu h_{0}m_{\bm{k}}=|\Delta_{0}s_{\bm{k}}|$, one of the bands reaches zero energy, namely, $E^{-}_{\bm{k}\nu}=0$, while $E^{+}_{\bm{k}\nu}\neq 0$. Thus, when the AM and SC have the same symmetry ($m_{\bm{k}}=s_{\bm{k}}$) a zero-energy flat band emerges as a bulk subgap state entirely induced by altermagnetism and akin to BFSs. As a result, the bulk density of states at zero energy becomes singular, similar to the divergent density of states at $E=\Delta_{0}$ for a conventional $s$-wave SC. This effect, and hence the flat band, can be detected via conductance, as we demonstrate in Fig. 2, where we plot the charge (spin) conductance for both $d_{x^{2}-y^{2}}$- and $d_{xy}$-wave AMs with $d_{x^{2}-y^{2}}$-wave superconductivity. The first feature to notice is that, at $h_{0}\neq\Delta_{0}$, the angle-resolved charge conductance $\bar{\sigma}^{e}(eV)$ is greatly enhanced at $eV=0$ as a result of the BFS ($E^{-}_{\bm{k}\nu}=0$) discussed above [Fig. 2(a)]. This gives rise to a zero-bias conductance peak (ZBCP) in the total charge conductance $\sigma_{\rm SC}^{e}(eV)$ at $h_{0}\neq\Delta_{0}$, which already appears at $Z=0$ but is enhanced as $Z$ enhances [Fig. 2(b)]. Thus, the ZBCP directly results is a bulk effect due to the BFS and it is not related to SABSs. By moving away from $eV=0$ and for $h_{0}\neq\Delta_{0}$, $\sigma_{\rm SC}^{e}(eV)$ develops a V-shape profile with spikes outside the gap [Fig. 2(c)], an effect that is similar to ferromagnetic SCs [39]. On the other hand, when $d_{xy}$-wave AM and $d_{x^{2}-y^{2}}$-wave SC coexist, no zero energy bulk flat band is found although the charge conductance still acquires finite subgap values following a U-shape profile for $h_{0}\neq 0$ but without any ZBCP, see Fig. 2(e,f).
To further unveil the consequences of the ZBCP and BFS found here, in Fig. 2(d) we show the spin conductance $\sigma_{\rm SC}^{s}(eV)$ found from Eqs. (6) as a function of $eV$ and distinct $h_{0}$ for a $d_{x^{2}-y^{2}}$-wave AM and a $d_{x^{2}-y^{2}}$-wave SC. For $h_{0}\neq 0$, a $\sigma_{\rm SC}^{s}(eV)\neq 0$ appears reflecting the role of $d_{x^{2}-y^{2}}$-wave altermagnetism, where $\Gamma_{\uparrow}^{\pm}\neq\Gamma_{\downarrow}^{\pm}$ [Eq. (5)]. We further see that the spin conductance bears multiple sign changes as $eV$ varies, and, interestingly, it exhibits an abrupt sign change for $h_{0}=\Delta_{0}$ at $eV=0$. Thus, the abrupt sign change in $\sigma_{\rm SC}^{s}(eV=0)$ is a direct consequence of the zero-energy bulk flat band when the symmetry of the AM and SC are the same. For distinct symmetries, as for a $d_{xy}$-wave AM and a $d_{x^{2}-y^{2}}$-wave SC in the inset of Fig. 2(f), the spin conductance vanishes due to absence of spin splitting for $d_{xy}$-wave AMs leading to $\Gamma_{\uparrow}^{\pm}=\Gamma_{\downarrow}^{\mp}$. Hence, having AMs and SCs with the same symmetry enables the realization of bulk zero-energy flat bands that can be detected via charge conductance and can be used to efficiently manipulate spin currents in an electric way without a net magnetization [102, 103, 104].
Figure 2: (a-d) Conductance for a junction having a $d_{x^{2}-y^{2}}$-SC with a $d_{x^{2}-y^{2}}$-AM (a-d), including the angle-resolved charge conductance at $Z=2$, $h_{0}=\Delta_{0}$ (a), the total zero-bias charge conductance as a function of $Z$ and $h_{0}$ (b), the total charge conductance as a function of eVeV for distinct $h_{0}$ at $Z=2$ (c), and the total spin conductance (d). (e,f) The same as in (a,c) for a $d_{xy}$-AM, with the inset in (f) its total spin conductance.
Surface curved and flat bands by $d_{xy}$-wave altermagnetism in $d_{xy}$-wave SCs.— It is well-known that $d_{xy}$-wave pair potentials lead to zero-energy surface flat states without AMs [12]. Here we are interested in analyzing how altermagnetism can control the surface states and how to induce them in regimes not present before. This is demonstrated in Fig. 3, where we plot the angle-resolved and total charge conductances. For a $d_{x^{2}-y^{2}}$-AM and $d_{xy}$-wave SC, the angle-resolved charge conductance $\bar{\sigma}^{e}$ as a function of $\theta$ shows that the surface states split in a curved shape, with the size of the splitting given by the anisotropic AM field, but remain degenerate at $\theta=\pi/4$ [Fig. 3(a)]. Due to this reason, the total charge conductance $\sigma_{\rm SC}^{e}$ gets suppressed as the AM field increases [Fig. 3(b,c)], going from a sharp ZBCP at $h_{0}=0$ into an almost constant subgap conductance for $h_{0}>0$ whose size determines the strength of altermagnetism. Fig. 3(c) further uncovers that this surface effect needs $Z>0$.
Figure 3: (a-c) Conductance for a junction having a $d_{xy}$-SC with a $d_{x^{2}-y^{2}}$-AM, including the angle-resolved charge conductance at $Z=2$, $h_{0}=\Delta_{0}$ (a), the total charge conductance as a function of eVeV for distinct $h_{0}$ at $Z=2$ (b), and the total zero-bias charge conductance as a function of $Z$ and $h_{0}$ (c). (d-f) The same as in (a-c) but for a $d_{xy}$-AM; the inset in (d) shows angle-resolved charge conductance at $h_{0}=1.5\Delta_{0}$.
For a $d_{xy}$-wave AM and $d_{xy}$-wave SC, where the superconducting gap nodes are intact, the surface flat bands inherent to $d_{xy}$-wave SCs remain robust against the AM field when $h_{0}<\Delta_{0}$ but are split for $h_{0}>\Delta_{0}$, see Fig. 3(d) and its inset. This leads to a ZBCP in $\sigma_{\rm SC}^{e}(eV)$ for $h_{0}<\Delta_{0}$, but it reduces when $h_{0}>\Delta_{0}$ into an almost constant value within $|eV|<h_{0}-\Delta_{0}$ [Fig. 3(e,f)]. Further insights are obtained by looking at the zero-bias angle-resolved charge conductance for the $d_{xy}$-AM,
$\bar{\sigma}_{\sigma}^{e}(eV=0)=1+\bigg{|}h_{0}/\Delta_{0}+\sqrt{h_{0}^{2}/\Delta_{0}^{2}-1}\bigg{|}^{-1}.$ (8)
For $h_{0}<\Delta_{0}$, the zero energy surface bound states emerge for any value of $\theta$ within the gap and thus give rise to $\bar{\sigma}_{\sigma}^{e}=2$ in Eq. (8). However, for $h_{0}>\Delta_{0}$, the generation of continuum states at zero energy causes $\bar{\sigma}_{\sigma}^{e}(0)$ to decrease. From this result, we can see that the ZBCP is influenced not only by the symmetry but also by the strength of $d_{xy}$-wave altermagnetism.
Figure 4: (a,b) Conductance for a junction having a chiral $p$-wave SC with a $d_{x^{2}-y^{2}}$-wave AM, including the angle-resolved charge conductance at $Z=2$, $h_{0}=0.5\Delta_{0}$ (a) and the total charge conductance as a function of eVeV for distinct $h_{0}$ (b). (c,d) The same quantities as in (a,b) but with a $d_{xy}$-AM.
Surface curved bands by $d$-wave altermagnetism in chiral $p$-wave SCs.— Chiral $p$-wave pairing was once believed to describe superconductivity in Sr2RuO4 with chiral edge states [105, 106, 107], and a recent study even supports that Sr2RuO4 supports intrinsic AM order [108]. Motivated by these ideas, here we consider a one-band model with chiral $p$-wave pairing and $d$-wave altermagnetism; in the End Matter, we also address $p$-wave pairing with $g$-wave altermagnetism. For a $d_{x^{2}-y^{2}}$-AM, we calculate the chiral SABSs by imposing $\Gamma_{\sigma}^{+}\Gamma_{\sigma}^{-}=-e^{2i\theta}$ in Eq. (7),
$E_{\text{ABS}\nu}=\Delta_{0}\sin\theta+\nu h_{0}\cos 2\theta\,.$ (9)
with $\nu=\pm$. This reveals that the altermagnet field $h_{0}$ induces a spin splitting in the chiral SABSs $E_{\text{ABS}}^{\pm}$, which appear with a curved dispersion with $\theta$ and also revealed by angle-resolved conductance in Fig. 4(a). The splitting arises due $\Gamma_{\uparrow}^{+}\Gamma_{\uparrow}^{-}\neq\Gamma_{\downarrow}^{+}\Gamma_{\downarrow}^{-}$ in a $d_{x^{2}-y^{2}}$-wave AM. At certain angles found from $\Delta_{0}\sin\theta=\pm h_{0}\cos 2\theta$, however, the SABSs remain degenerate at zero energy [Fig. 4(a)], which affects their dispersion and makes them to acquire a quasi-flat band profile for a range of $\theta$. These quasi-flat bands then induce spikes in the total charge conductance Fig. 4(b) whose separation for $h_{0}>\Delta_{0}/(2\sqrt{2})$ is given by $\Delta_{0}^{2}/\left(4h_{0}\right)+2h_{0}$ from $\partial E_{\text{ABS}\nu}/\partial\theta=0$ and hence determines the strength of altermagnetism.
When $d_{xy}$-wave altermagnetism is present in chiral $p$-wave SCs, we obtain the chiral SABSs to be given by
$E_{\text{ABS}\nu}=\sin\theta\sqrt{\Delta_{0}^{2}-4h_{0}^{2}\sin^{2}\theta}\,$, (10)
which shows that SABSs acquire a curved dispersion with $\theta$ but are not split in the presence of $d_{xy}$-AM because here $\Gamma_{\uparrow}^{+}\Gamma_{\uparrow}^{-}=\Gamma_{\downarrow}^{+}\Gamma_{\downarrow}^{-}$. Moreover, Eq. (10) further unveils that the observation of such curved SABSs require high angles, $|\theta|>\arcsin^{-1}[\Delta_{0}/\left(2h_{0}\right)]$, as demonstrated by the angle-resolved charge conductance in Fig. 4(c). The corresponding total charge conductance then exhibits a broadened ZBCP for $|h_{0}|<\Delta_{0}$ which then splits around $eV=0$ for $|h_{0}|>\Delta_{0}$, as shown in Fig. 4(d).
Figure 5: (a-c) Conductance for a junction having a chiral $d$-wave SC with a $d_{x^{2}-y^{2}}$-wave AM, including the angle-resolved charge conductance at $Z=2$, $h_{0}=\Delta_{0}$ (a), the total charge conductance as a function of eVeV for distinct $h_{0}$ (b), and the zero-bias total charge conductance as a function of $Z$ and $h_{0}$. (d-f) The same quantities as in (a-c) but with a $d_{xy}$-AM
Surface flat bands by $d_{x^{2}-y^{2}}$-wave altermagnetism in chiral $d$-wave SCs.— To have a complete picture of altermagnetism inducing subgap states, here we address it when chiral $d$-wave pairing coexists with the two types of AMs. In the absence of AMs, it is well-known that chiral $d$-wave SCs host 2 chiral SABSs [109, 110, 111, 112]. Thus, under $d_{x^{2}-y^{2}}\,$-AM, it is reasonable to expect that such SABSs are split into 4 by $d_{x^{2}-y^{2}}\,$-AM. Actually, the obtained SABSs are given by
$E_{\text{ABS}\nu}=-\left(\Delta_{0}+\nu h_{0}\right)\text{sgn}\left(\theta\right)\cos 2\theta,$ (11)
where $\nu=\pm$ and $\text{sgn}(.)$ denotes the sign operation. The split SABSs can be seen from Eq. (11). Interestingly, we verified that the dispersion of one SABS becomes steeper, while the other flattens by increasing $h_{0}$ and gets completely flat at zero energy when $h_{0}=\Delta_{0}$ for all $k_{y}$, as shown in the angle-resolved charge conductance in Fig. 5(a). The zero-energy flat band further leads to a ZBCP in the total charge conductance at $h_{0}=\Delta_{0}$ [Fig. 5(b)], while the subgap conductance becomes less prominent for $h_{0}\neq\Delta_{0}$ Fig. 5(b,c). For the $d_{xy}$-AM, the pair of SABSs acquire instead a curved dispersion due to altermagnetism which gives rise to a small enhancement of the subgap conductance, as seen in Figs.5(d,e). Moreover, such SABSs do not flatten as the $d_{xy}$-AM field enhances, which implies that the zero-biased conductance reduces by increasing $Z$ [Fig. 5(f)].
Summary.— In this work, we have demonstrated rich properties of subgap bands and surface bound states where superconductivity and altermagnetism coexist. It shows that the zero-energy states can be enhanced when the symmetry of altermagnetism and pairings are the same, leading to a zero-biased conductance peak without any formation of surface bound states. We verified the robustness of the flat surface Andreev bound states arising from unconventional nodal pairings, which can be seriously influenced by either the symmetry or the magnitude of altermagnetism. It is also found that in chiral superconductors, the chiral Andreev bound states can be split, curved, or flattened by altermagnetism. In terms of experimental test of our predictions, it might be possible exploit the reported altermagnets, such as KV2Se2 [113] and various unconventional superconductors, such as cuprates [114]. Moreover, recent studies have shown that both the magnetic strength and orientation of altermagnets are electrically controllable [115, 116], providing an accessible way to change altermagnetic fields. For future applications, the subgap states are shown to bear large spin conductance with zero net magnetic field, useful for making functional superconducting spintronic devices.
Acknowledgements.— B. L. acknowledges financial support from the National Natural Science Foundation of China (project 12474049). P.M. and P.B. acknowledge support by the Spanish CM “Talento Program” project No. 2019-T1/IND-14088 and No. 2023-5A/IND-28927, the Agencia Estatal de Investigación project No. PID2020-117992GA-I00 and No. CNS2022-135950 and through the “María de Maeztu” Programme for Units of Excellence in R&D (CEX2023-001316-M). J. C. acknowledges financial support from the Swedish Research Council (Vetenskapsrådet Grant No. 2021-04121) and the Carl Trygger’s Foundation (Grant No. 22: 2093). Y. T.  acknowledges financial support from JSPS with Grants-in-Aid for Scientific Research (KAKENHI Grants Nos. 23K17668, 24K00583, 24K00556, 24K00578, 25H00609 and 25H00613). Y. F. acknowledges financial support from the Sumitomo Foundation and numerical calculation support from Okayama University.