### v4

# Engineering subgap states in superconductors by the symmetry of altermagnetism

Abstract
Combining superconducting and magnetic materials is a promising path to generate exotic interface subgap states. In this regard, altermagnetism is particularly interesting because it lifts spin degeneracy while providing tailored anisotropy of spin splittings. Here, we investigate the realization and control of subgap states by using the symmetry contrast between altermagnetic fields and unconventional pairings. When the symmetries of altermagnetism and unconventional superconductivity align, we demonstrate the emergence of bulk zero-energy flat bands as the Bogoliubov Fermi surface, giving rise to a zero-bias conductance peak. The symmetry and strength of $d$-wave altermagnets strongly affect the surface Andreev states from $d$-wave and chiral $d$- and $p$-wave superconductors. As a result, distinct types of subgap states are realized, including curved and flat bands, that can be detected by tunneling spectroscopy. Our results offer a solid route for designing and manipulating subgap states in superconducting systems, which can be useful for functionalizing superconducting devices.
Introduction.— The ability to induce and control subgap states in superconducting systems is a central topic in condensed matter physics, with a profound fundamental relevance and high potential for quantum applications. One example of such subgap states is the surface Andreev bound states (SABSs) [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13], which emerge due to the anisotropic nature of the superconducting pair potentials and can be detected by tunneling spectroscopy [6]. Interestingly, SABSs in unconventional superconductors (SCs) can lead to Majorana states [14, 15, 16, 12], which hold promise for designing topological quantum bits [17, 18, 19, 20, 21, 22]. Moreover, certain superconducting setups can also host bulk subgap states, such as Bogoliubov-Fermi surfaces (BFSs) [23, 24, 25, 26, 27, 28, 29, 30, 31, 32], which signal the nontrivial topological nature of superconductivity [33, 27]. For practical purposes, subgap states need to be manipulated, often carried out by coupling external fields to their intrinsic quantum numbers [34, 35, 36], such as orbital or spin. Although this can indeed be achieved by ferromagnets [37, 38, 39], their unavoidable finite magnetization and stray fields make them detrimental for superconductivity [40], and therefore, other materials are required.
Figure 1: (a) An altermagnet (AM, orange) is placed on top of a superconductor (SC, green) to engineer subgap states. A normal metal probe (N, blue) is coupled to SC at $x=0$ near the AM-SC interface, achieved by adjusting the thickness of the insulating substrate (grey) beneath N. (b) Illustrations of the selected AM orderings in momentum space where the red and blue colors indicate up and down spins. (c) Table listing the type of emergent subgap states for several AM-SC combinations. Labels “flat”, “split”, and “curved” indicate when the dispersion of subgap states is respectively zero, spin-split, or drastically varying.
With the recent advent of altermagnetism [41, 42], a powerful avenue has emerged for coupling spins in superconductors thanks to the altermagnetic spin-orbit properties [43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53]. Altermagnets (AMs) are spin-compensated magnetic materials with a nonrelativistic momentum-dependent spin splitting and anisotropic spin-polarized Fermi surfaces [54], which result in zero net magnetization and hence vanishing stray fields unlike ferromagnets [55, 56]. In combination with SCs, AMs have already been shown to host intriguing properties [53], such as orientation-dependent Andreev reflections [57, 58, 59, 60, 61, 62], anomalous Josephson currents [63, 64, 65, 66, 67, 68, 69], nontrivial light-matter coupling [70, 71], superconducting diodes [72, 73, 74], spin-triplet Cooper pairs [75, 76, 77, 78, 79], Majorana states [80, 81, 82, 83, 84, 85], among other phenomena [86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96]. Despite the impressive advances, the use of altermagnetism for realizing subgap states is still an open problem.
In this work, we demonstrate the emergence and control of highly tunable subgap states by using altermagnetism in unconventional SCs. To show this, we focus on a hybrid system composed of a two-dimensional SC with altermagnetism due to its proximity to an AM [Fig. 1(a)], where a normal probe enables the detection of bulk and surface subgap states. We consider various types of unconventional SCs combined with $d$-wave AMs [Fig. 1(b)], which allows us to reveal different types of subgap states, including bulk zero energy flat bands and SABSs, as listed in Fig. 1(c). Such subgap states are driven by the symmetry contrast between competing superconducting and magnetism orders and give rise to strong charge conductance signatures, enabling their detection and also the identification of the type of altermagnetism supporting them.
Model.— We aim to harness altermagnetism to realize and detect subgap states in unconventional SCs. We thus consider a bilayer AM-SC device where the coupling between AM and SC accounts for a proximity-induced AM field in the SC. Then, a two-dimensional normal metal lead N is connected to the SC at $x=0$ to probe the properties of altermagnetic superconductors, see Fig. 1(a). The N-SC planar junction is modeled by $\mathcal{H}=(1/2)\sum\nolimits_{\bm{k}}\hat{c}_{\bm{k}}^{{\dagger}}\mathcal{H}_{\bm{k}}\hat{c}_{\bm{k}}$, where the Nambu spinor is given by $\hat{c}_{\bm{k}}=(c_{\bm{k}\uparrow},c_{\bm{k}\downarrow},c_{-\bm{k}\uparrow}^{{\dagger}},c_{-\bm{k}\downarrow}^{{\dagger}})^{T}$, $\bm{k}=(k_{x},k_{y})$, and
$\mathcal{H}_{\bm{k}}=\left[\xi_{\bm{k}}+U(x)\right]\hat{\tau}_{z}+\mathcal{M}_{\bm{k}}(x)\hat{s}_{z}\hat{\tau}_{z}+\Delta_{\bm{k}}(x).$ (1)
Here, $\xi_{\bm{k}}=\hbar^{2}|\bm{k}|^{2}/(2m)-\mu$ is the kinetic energy, with $\mu$ being the chemical potential, $U(x)=U_{0}\delta_{x}$ is the interface barrier potential at $x=0$, with $U_{0}$ the barrier strength defining the parameter $Z=mU_{0}/(\hbar^{2}k_{F})$ that characterizes the tunneling ($Z\gg 1$) and full transparent ($Z=0$) regimes, $k_{F}=\sqrt{2m\mu/\hbar^{2}}$, and $\hat{s}_{i}(\hat{\tau}_{i})$ is the $i$-th Pauli matrix in spin (Nambu) space. Moreover, $\mathcal{M}_{\bm{k}}(x)=\Theta(x)h_{0}m_{\bm{k}}$ is the altermagnetic field, with $h_{0}$ denoting the altermagnetic strength and $m_{\bm{k}}$ the anisotropic factor, and, without loss of generality, the Neel vector of the AM is chosen along the $z$-axis. The unconventional superconducting pair potential $\Delta_{\bm{k}}(x)=\Theta(x)\hat{\Delta}_{\bm{k}}$ is only finite in the SC. We assumed $h_{0}\ll\mu$ so that the wavevectors at low energies can be approximated by $\bm{k}=k_{F}\bm{\hat{k}}$, with the unit vector $\bm{\hat{k}}=(\hat{k}_{x},\hat{k}_{y})$.
We consider highly relevant AMs exhibiting $d$-wave parity [Fig. 1(b)] expected in, e.g., RuO2 and KV2Se2 [42, 97], and modeled by specific anisotropic factors $m_{\bm{k}}$. We have $m_{\bm{k}}=(\hat{k}_{x}^{2}-\hat{k}_{y}^{2})\cos 2\alpha+2\hat{k}_{x}\hat{k}_{y}\sin 2\alpha$, with $\alpha=0$ for $d_{x^{2}-y^{2}}$-wave AMs and $\alpha=\pi/4$ for $d_{xy}$-wave AMs. For the unconventional SCs, we consider four types of experimentally relevant pairings [98, 12]. For $d$-wave SCs, we have $\hat{\Delta}_{\bm{k}}=-\Delta_{0}s_{\bm{k}}\hat{s}_{y}\hat{\tau}_{y}$, with $s_{\bm{k}}=(\hat{k}_{x}^{2}-\hat{k}_{y}^{2})\cos 2\chi+2\hat{k}_{x}\hat{k}_{y}\sin 2\chi$. Here, i) $\chi=0$ represents $d_{x^{2}-y^{2}}$-wave pairing and ii) $\chi=\pi/4$ $d_{xy}$-wave. These two pairing states are related by a rotation with respect to the orientation of the interface with the metallic probe. Chiral SCs display iii) chiral $p$-wave pairing $\hat{\Delta}_{\bm{k}}=\Delta_{0}\hat{s}_{x}(\hat{k}_{x}\hat{\tau}_{x}-\hat{k}_{y}\hat{\tau}_{y})s_{\bm{k}}$ or iv) chiral $d$-wave pairing $\hat{\Delta}_{\bm{k}}=-\Delta_{0}(\hat{k}_{x}^{2}-\hat{k}_{y}^{2})\hat{s}_{y}\hat{\tau}_{y}s_{\bm{k}}-2\Delta_{0}\hat{k}_{x}\hat{k}_{y}\hat{s}_{y}\hat{\tau}_{x}s_{\bm{k}}$. In both cases $s_{\bm{k}}=1$. Here, $\Delta_{0}$ is the gap amplitude in the presence of the AM field and can be suppressed [92, 99].
The quasiparticle energy dispersion in the SC with altermagnetism is given by
$E_{\bm{k}\nu}^{\pm}=\nu h_{0}m_{\bm{k}}\pm\sqrt{\xi_{\bm{k}}^{2}+|\Delta_{0}s_{\bm{k}}|^{2}}\,$, (2)
where $\nu=\pm$ indicating two possible anisotropic energy shifts ($\pm h_{0}m_{\bm{k}}$) of subbands. From Eq. (2), it is easy to see that, for $\nu h_{0}m_{\bm{k}}=|\Delta_{0}s_{\bm{k}}|$, we get $E^{+}_{\bm{k}\nu}\neq 0$ and $E^{-}_{\bm{k}\nu}=0$, the latter signaling a zero energy bulk flat band akin to BFSs [23]. This is an example of a bulk subgap state induced by altermagnetism, and its detection can be carried out via conductance in the setup of Fig. 1(a). Notably, this setup also allows us to search for altermagnetism-induced subgap states at the surface, which can be directly accessed via conductance as we explain below.
To obtain the conductance, we employ the scattering approach, which is based on the wavefunctions of injected electrons with spin $\sigma=\uparrow,\downarrow$. Due to translational invariance along $y$, the wavefunctions are $\Psi_{\sigma}(x,k_{y})=\psi_{\sigma}(x)e^{ik_{y}y}$. In the normal side, $\psi_{\sigma}(x)$ is given by
$\psi_{\sigma}(x<0)=[\hat{A}_{\sigma}+a_{\sigma}\hat{B}_{\sigma}]e^{ik_{x}x}+b_{\sigma}\hat{A}_{\sigma}e^{-ik_{x}x}\,$, (3)
where $\hat{A}_{\uparrow}=\left(1,0,0,0\right)^{T}$, $\hat{A}_{\downarrow}=\left(0,1,0,0\right)^{T}$, $\hat{B}_{\downarrow}=\left(0,0,1,0\right)^{T}$ and $\hat{B}_{\uparrow}=\left(0,0,0,1\right)^{T}$, $k_{x}=k_{F}\cos\theta$, and $k_{y}=k_{F}\sin\theta$. The wave functions in the SC can be written in a similar way. Then, the coefficients $a_{\sigma}$ and $b_{\sigma}$ are obtained from the boundary conditions $\psi(0^{+})=\psi(0^{-})$, $\partial_{x}\psi(0^{+})-\partial_{x}\psi(0^{-})=(2mU_{0}/\hbar^{2})\psi(0)$. The angle-resolved charge conductance is found as [100] $\bar{\sigma}_{\sigma}^{e}=1+|a_{\sigma}|^{2}-|b_{\sigma}|^{2}$, which gives
$\bar{\sigma}_{\sigma}^{e}=\bar{\sigma}_{n}\frac{1+\bar{\sigma}_{n}|\Gamma_{\sigma}^{+}|^{2}+(\bar{\sigma}_{n}-1)|\Gamma_{\sigma}^{+}\Gamma_{\sigma}^{-}|^{2}}{|\Lambda-\eta(1-\bar{\sigma}_{n})\Gamma_{\sigma}^{+}\Gamma_{\sigma}^{-}|^{2}}\,$, (4)
with $\bar{\sigma}_{n}=1/[1+Z^{2}/(\cos^{2}\theta)]$. Here, $\Gamma_{\sigma}^{\pm}$ are given by
$\Gamma_{\sigma}^{\pm}=\frac{\Delta_{0}s_{\bm{k}_{\pm}}}{E_{\sigma,\bm{k}_{\pm}}+\sqrt{E_{\sigma,\bm{k}_{\pm}}^{2}-\Delta_{0}^{2}s_{\bm{k}_{\pm}}^{2}}}.$ (5)
where $E_{\uparrow,\bm{k}_{\pm}}=E-\mathcal{M}_{\bm{k}_{\pm}}$ and $E_{\downarrow,\bm{k}_{\pm}}=E+\mathcal{M}_{\bm{k}_{\pm}}$, with $\bm{k}_{+}$, $\bm{k}_{-}$ are given by $\bm{k}_{\pm}=k_{F}\left(\pm\cos\theta,\sin\theta\right)$. Moreover, $\Lambda=1$ for $d_{x^{2}-y^{2}}$- and $d_{xy}$-wave SCs; $\Lambda=e^{2i\theta}$ for chiral $p$-wave SCs; $\Lambda=e^{4i\theta}$ for chiral $d$-wave SCs. Also, $\eta=1$ for spin-singlet SCs (cases i,ii,iv), while $\eta=-1$ for spin-triplet SCs (case iii). Hence, the charge conductance in the superconducting state is calculated as
$\sigma_{\mathrm{SC}}^{e}=\frac{e^{2}}{h}\int_{-\frac{\pi}{2}}^{\frac{\pi}{2}}\bar{\sigma}^{e}\cos\theta d\theta,$ (6)
where $\bar{\sigma}^{e}=\bar{\sigma}_{\uparrow}^{e}+\bar{\sigma}_{\downarrow}^{e}$ is the angle-resolved charge conductance. Later we normalize the conductance by $\sigma_{N}=(2e^{2}/h)\int\bar{\sigma}_{n}\cos\theta d\theta$.
At this point, it is important to highlight that, to access the surface ABSs states via conductance, the necessary condition is to set $\bar{\sigma}_{n}\rightarrow 0$ and find the zeroes of the denominator in Eq. (4), which leads to
$\Gamma_{\sigma}^{+}\Gamma_{\sigma}^{-}=\Lambda/\eta\,.$ (7)
This condition corresponds to the poles of the associated scattering matrix. Thus, the surface ABSs are obtained from Eq. (7). In what follows, we explore how the bulk subgap state discussed below Eq. (2) as well as surface ABSs can be induced by altermagnetism and detected by conductance in unconventional SCs. Our results in the continuum model are the same with the lattice model shown in the Supplementary Material [99].
Bulk and surface flat bands by $d$-wave AMs in $d$-wave SCs.— We begin by analyzing the junction formed by $d_{x^{2}-y^{2}}$-wave SC with $d$-wave altermagnetism. Before delving into the conductance, we remind our discussion below Eq. (2), where we show that at $\nu h_{0}m_{\bm{k}}=|\Delta_{0}s_{\bm{k}}|$, one of the bands reaches zero energy, namely, $E_{\bm{k}\nu}^{-}=0$, while $E_{\bm{k}\nu}^{+}\neq 0$. Thus, when the AM and SC have the same symmetry ($m_{\bm{k}}=s_{\bm{k}}$) a zero-energy flat band emerges as a bulk subgap state entirely induced by altermagnetism and akin to BFSs. Indeed, this flat band is only possible due to the accidental disappearance of a gaped spectrum at $E=0$ while the superconducting gap still exists at $|E|<2\Delta_{0}$, maintaining superconducting coherence. As a result, the bulk density of states at zero energy becomes singular, similar to the divergent density of states at $E=\Delta_{0}$ for a conventional $s$-wave SC. This effect, and hence the flat band, can be detected via conductance, as we demonstrate in Fig. 2, where we plot the charge conductance showing nontrivial subgap states with $d$-wave AMs and $d$-wave SCs. The first feature to notice is that, for a $d_{x^{2}-y^{2}}$-wave SC with $d_{x^{2}-y^{2}}$-wave AM, the angle-resolved charge conductance $\bar{\sigma}^{e}(eV)$ at $h_{0}=\Delta_{0}$ is greatly enhanced at $eV=0$ as a result of the BFS ($E_{\bm{k}\nu}^{-}=0$) discussed above [Fig. 2(a)]. This enhancement gives rise to a zero-bias conductance peak (ZBCP) in the total charge conductance $\sigma_{\mathrm{SC}}^{e}(eV)$ at $h_{0}=\Delta_{0}$ [Fig. 2(b)]. The ZBCP stems from bulk BFSs and thus its width is very narrow [101]. By moving away from $eV=0$ and for $h_{0}\neq\Delta_{0}$, $\sigma_{\mathrm{SC}}^{e}(eV)$ develops a V-shape profile with spikes outside the gap, an effect that is similar to ferromagnetic SCs [39]. It is worth mentioning that when $d_{xy}$-wave AM and $d_{x^{2}-y^{2}}$-wave SC coexist, no zero energy bulk flat band can be found according to Eq. (2).
Figure 2: Conductance for a junction with various combinations of $d$-wave SCs and $d$-wave AMs. The AM is $d_{x^{2}-y^{2}}$-wave in panels (a-c) and $d_{xy}$-wave in (d). SC pairing is $d_{x^{2}-y^{2}}$-wave in (a), (b) and $d_{xy}$-wave in (c), (d). Panel (a) show angle-resolved conductance at $h_{0}=\Delta_{0}$. For all panels $Z=2$.
Now we turn to $d_{xy}$-wave SCs. It is well-known that $d_{xy}$-wave pair potentials lead to zero-energy surface flat states without AMs [12]. Here we are interested in analyzing how AM can control the surface states and how to induce them in regimes not previously present. This is demonstrated in Figs. 2(c,d), where we plot the total charge conductance. For a $d_{x^{2}-y^{2}}$-wave AM and $d_{xy}$-wave SC, the total charge conductance $\sigma_{\mathrm{SC}}^{e}$ gets suppressed as the AM field increases, going from a sharp ZBCP at $h_{0}=0$ into an almost constant subgap conductance for $h_{0}>0$ whose size determines the strength of altermagnetism [Fig. 2(c)]. For a $d_{xy}$-wave AM and $d_{xy}$-wave SC, where the superconducting gap nodes are intact, the surface flat bands inherent to $d_{xy}$-wave SCs remain robust against the AM field when $h_{0}<\Delta_{0}$, see Fig. 2(d). This leads to a ZBCP in $\sigma_{\mathrm{SC}}^{e}(eV)$ for $h_{0}<\Delta_{0}$, but it reduces when $h_{0}>\Delta_{0}$ into an almost constant value within $|eV|<h_{0}-\Delta_{0}$. Further insights are obtained by looking at the zero-bias angle-resolved charge conductance for the $d_{xy}$-wave AM,
$\bar{\sigma}_{\sigma}^{e}(eV=0)=1+\bigg|h_{0}/\Delta_{0}+\sqrt{h_{0}^{2}/\Delta_{0}^{2}-1}\bigg|^{-1}.$ (8)
For $h_{0}<\Delta_{0}$, the zero energy surface bound states emerge for any value of $\theta$ within the gap and thus give rise to $\bar{\sigma}_{\sigma}^{e}=2$ in Eq. (8). However, for $h_{0}>\Delta_{0}$, the generation of continuum states at zero energy causes $\bar{\sigma}_{\sigma}^{e}(0)$ to decrease. From this result, we can see that the ZBCP is influenced not only by the symmetry but also by the strength of $d_{xy}$-wave altermagnetism.
Figure 3: Conductance for junctions with chiral $p$-wave SCs. AM is $d_{x^{2}-y^{2}}$ in (a)(c) and $d_{xy}$ in (b)(d). (a)(b) angle-resolved charge conductance at $h_{0}=\Delta_{0}$. (c)(d) total conductance for distinct values of $h_{0}$. $Z=2$ for all panels.
Surface curved bands by $d$-wave altermagnetism in chiral $p$-wave SCs.— Chiral $p$-wave pairing was once believed to describe superconductivity in Sr2RuO4 with chiral edge states [102, 103, 104], and a recent study even supports that Sr2RuO4 supports intrinsic AM order [105]. Motivated by these ideas, here we consider a one-band model with chiral $p$-wave pairing and $d$-wave altermagnetism; in the End Matter, we also address $p$-wave pairing with $g$-wave altermagnetism. For a $d_{x^{2}-y^{2}}$-wave AM, we calculate the chiral SABSs by imposing $\Gamma_{\sigma}^{+}\Gamma_{\sigma}^{-}=-e^{2i\theta}$ in Eq. (7),
$E_{\text{ABS}\nu}=\Delta_{0}\sin\theta+\nu h_{0}\cos 2\theta,$ (9)
with $\nu=\pm$. This reveals that the altermagnet field $h_{0}$ induces a spin splitting in the chiral SABSs $E_{\text{ABS}}^{\pm}$, which appear with a curved dispersion with $\theta$ and also revealed by angle-resolved conductance in Fig. 3(a). The splitting arises due $\Gamma_{\uparrow}^{+}\Gamma_{\uparrow}^{-}\neq\Gamma_{\downarrow}^{+}\Gamma_{\downarrow}^{-}$ in a $d_{x^{2}-y^{2}}$-wave AM and there are thus two species of SABSs as shown in Fig. 3(a). At certain angles found from $\Delta_{0}\sin\theta=\pm h_{0}\cos 2\theta$, however, the SABSs remain degenerate at zero energy [Fig. 3(a)], which affects their dispersion and makes them to acquire a quasi-flat band profile for a range of $\theta$. These quasi-flat bands then induce spikes in the total charge conductance whose separation for $h_{0}>\Delta_{0}/(2\sqrt{2})$ [Fig. 3(c)] is given by $\Delta_{0}^{2}/\left(4h_{0}\right)+2h_{0}$ from $\partial E_{\text{ABS}\nu}/\partial\theta=0$ and hence determines the strength of altermagnetism.
When $d_{xy}$-wave altermagnetism is present in chiral $p$-wave SCs, we obtain the chiral SABSs to be given by
$E_{\text{ABS}\nu}=\sin\theta\sqrt{\Delta_{0}^{2}-4h_{0}^{2}\sin^{2}\theta}\,$, (10)
which shows that SABSs acquire a curved dispersion with $\theta$ but are not split in the presence of $d_{xy}$-wave AM because here $\Gamma_{\uparrow}^{+}\Gamma_{\uparrow}^{-}=\Gamma_{\downarrow}^{+}\Gamma_{\downarrow}^{-}$. Moreover, Eq. (10) further unveils that the observation of such curved SABSs require $|\theta|<\arcsin^{-1}[\Delta_{0}/\left(2h_{0}\right)]$, as demonstrated by the angle-resolved charge conductance in Fig. 3(b). This curved dispersion of SABSs can result in a quasi-flat band at large angle and the corresponding total charge conductance then exhibits a broadened ZBCP for $|h_{0}|<\Delta_{0}$ which then splits around $eV=0$ for $|h_{0}|>\Delta_{0}$, as shown in Fig. 3(d).
Figure 4: Conductance for junctions with chiral $d$-wave SCs with a $d$-wave AM. (a) angle-resolved charge conductance at $h_{0}=\Delta_{0}$ and (b) total conductance for distinct values of $h_{0}$ for a $d_{x^{2}-y^{2}}$-wave AM. (c) zero bias conductance vs $\alpha$ and $h_{0}$ where $\alpha$ denotes the crystal orientation of AM. (d) Same junction with (c) but the total conductance as a function of $eV$ and $\alpha$ at $h_{0}=\Delta_{0}$. $Z=2$ for all panels.
Surface flat bands by $d_{x^{2}-y^{2}}$-wave altermagnetism in chiral $d$-wave SCs.— To have a complete picture of altermagnetism inducing subgap states, here we address it when chiral $d$-wave pairing coexists with $d$-wave AMs. In the absence of AMs, it is well-known that chiral $d$-wave SCs host 2 chiral SABSs [106, 107, 108, 109]. Thus, under $d_{x^{2}-y^{2}}\,$-AM, it is reasonable to expect that such SABSs are split into 4 by $d_{x^{2}-y^{2}}\,$-AM. Actually, the obtained SABSs are given by
$E_{\text{ABS}\nu}=-\left(\Delta_{0}+\nu h_{0}\right)\text{sgn}\left(\theta\right)\cos 2\theta,$ (11)
where $\nu=\pm$ and $\text{sgn}(.)$ denotes the sign operation. The split SABSs can be seen from Eq. (11). Interestingly, we verified that the dispersion of one SABS becomes steeper, while the other flattens by increasing $h_{0}$ and gets completely flat at zero energy when $h_{0}=\Delta_{0}$ for all $k_{y}$, as shown in the angle-resolved charge conductance in Fig. 4(a). The zero-energy flat band further leads to a ZBCP in the total charge conductance at $h_{0}=\Delta_{0}$ [Fig. 4(b)], while the subgap conductance becomes less prominent for $h_{0}\neq\Delta_{0}$. For the $d_{xy}$-AM, the pair of SABSs have no spin-splittings but acquire instead a curved dispersion due to altermagnetism. It is noted that the emergence of ZBCP can be robust against a certain twist between AM and SC. As Figs.4(c,d) show, there is a conductance peak at $h_{0}=\Delta_{0}$ for a large range around $\alpha=\pi/2$ ($d_{x^{2}-y^{2}}$-AM) and disapear at $\alpha=3\pi/4$ ($d_{xy}$-AM). We thus view it as the resilient feature for experimental detection of the flat SABSs in chiral $d$-wave SCs.
Summary.— In this work, we have demonstrated rich properties of subgap bands and surface bound states where superconductivity and altermagnetism coexist. It shows that the zero-energy states can be enhanced when the symmetry of altermagnetism and pairings are the same, leading to a zero-biased conductance peak without any formation of surface bound states. We verified the robustness of the flat surface Andreev bound states arising from unconventional nodal pairings, which can be seriously influenced by either the symmetry or the magnitude of altermagnetism. It is also found that in chiral superconductors, the chiral Andreev bound states can be split, curved, or flattened by altermagnetism. Our study reveals the importance of the anisotropic orbital symmetry as well as the lifting of spin degeneracy by altermagnetism in the formation of subgap states. Notably, the integration of altermagnet-superconductor heterostructures is expected to induce spin-triplet pairings [110] and the characteristics of resulting subgap states by spin-orbit and impurity effect would be an interesting topic for future research. In terms of experimental test of our predictions, it might be possible to exploit the reported altermagnets, such as KV2Se2 [97] and various unconventional superconductors, such as cuprates [111]. Moreover, recent studies have shown that both the magnetic strength and orientation of altermagnets are electrically controllable [112, 113], providing an accessible way to change altermagnetic fields. It has also been proposed that chemical doping has the potential to control the AM field [114]. For future applications, the subgap states are shown to modulate conductance spectra in the absence of a net magnetic field, useful for making field-free functional superconducting devices.
Acknowledgements.— B. L. thanks F. M. Qu for useful discussions related to this work and acknowledges financial support from the National Natural Science Foundation of China (project 12474049) and Beijing National Laboratory for Condensed Matter Physics (2025BNLCMPKF011). P.M. and P.B. acknowledge support by the Spanish CM “Talento Program” project No. 2019-T1/IND-14088 and No. 2023-5A/IND-28927, the Agencia Estatal de Investigación project No. PID2020-117992GA-I00, No. PID2024-157821NB-I00 and No. CNS2022-135950 and through the “María de Maeztu” Programme for Units of Excellence in R&D (CEX2023-001316-M). J. C. acknowledges financial support from the Swedish Research Council (Vetenskapsrådet Grant No. 2021-04121) and the Carl Trygger’s Foundation (Grant No. 22: 2093). Y. T. acknowledges financial support from JSPS with Grants-in-Aid for Scientific Research (KAKENHI Grants Nos. 23K17668, 24K00583, 24K00556, 24K00578, 25H00609 and 25H00613). Y. F. acknowledges financial support from the Sumitomo Foundation and numerical calculation support from Okayama University.